# OpenRLHF: An Easy-to-use, Scalable and High-performance RLHF Framework

### v5

@@ Abstract
Large Language Models (LLMs) fine-tuned via Reinforcement Learning from Human Feedback (RLHF) and Reinforcement Learning with Verifiable Rewards (RLVR) significantly improve the alignment of human-AI values and further raise the upper bound of AI capabilities, particularly in reasoning-intensive, long-context Chain-of-Thought (long-CoT) tasks. However, existing RLHF (or RLVR) frameworks commonly face challenges such as inference bottlenecks and complexity barriers, restricting their accessibility for newcomers. To bridge this gap, we introduce OpenRLHF, a user-friendly, scalable, and easy-to-learn open-source RLHF framework built upon Ray, vLLM, DeepSpeed, and HuggingFace Transformers, featuring a simplified design, clear code structure, and comprehensive documentation to facilitate entry for researchers and practitioners. Experimental results show that OpenRLHF achieves superior training efficiency with speedups ranging from 1.22× to 1.68× across different model sizes compared to state-of-the-art frameworks, while requiring significantly fewer lines of code for implementation. OpenRLHF is publicly available at https://github.com/OpenRLHF/OpenRLHF, and has already been adopted by leading institutions to accelerate RLHF research and learning.
OpenRLHF: An Easy-to-use, Scalable and High-performance RLHF Framework
Jian Hu, Xibin Wu, Wei Shen, Jason Klein Liu, Zilin Zhu Weixun Wang, Songlin Jiang, Haoran Wang, Hao Chen, Bin Chen, Weikai Fang, Xianyu, Yu Cao, Haotian Xu, Yiming Liuthanks: Complete contributor list available in Appendix C. [Uncaptioned image] Team janhu9527@gmail.com
@@ 1 Introduction
Large Language Models (LLMs) fine-tuned via Reinforcement Learning from Human Feedback (RLHF) and Reinforcement Learning with Verifiable Rewards (RLVR) have markedly advanced human-AI alignment and elevated the upper bound of AI capabilities Christiano et al. (2017); Stiennon et al. (2020); Guo et al. (2025); Shen et al. (2025). These approaches enable models to better conform to human intentions and values while achieving superior reasoning performance. Notably, models such as GPT-4 Achiam et al. (2023), DeepSeek-R1 Guo et al. (2025), and Claude Askell et al. (2021) excel at complex reasoning tasks by generating detailed step-by-step rationales, commonly referred to as Chain-of-Thought (CoT) outputs.
However, RLHF and RLVR training methodologies—especially those employing Proximal Policy Optimization (PPO)—face significant computational challenges. In particular, the inference phase often accounts for over 90% of the total RLHF (or RLVR) runtime, as models need to generate thousands of tokens during each inference step. Consequently, there is an increasing demand for efficient and scalable frameworks that reduce inference overhead and simplify distributed RLHF and RLVR training workflows.
Existing RLHF and RLVR systems such as DeepSpeed-Chat Yao et al. (2023), TRL von Werra et al. (2020), and ColossalChat Li et al. (2023) have made noteworthy progress in distributed computation and memory efficiency. Conversely, industrial-grade solutions like Nemo-aligner Shen et al. (2024a) and ChatLearn alibaba (2017) offer advanced optimizations at the modeling and framework levels. While Verl Sheng et al. (2024), a framework proposed after our initial development, also provides sophisticated optimizations (e.g., its 3D-Hybrid engine), these industrial solutions generally feature tightly coupled and specialized designs that introduce considerable complexity, steep learning curves, and accessibility barriers for newcomers and academic researchers. Yet, their tightly coupled and specialized designs introduce considerable complexity, steep learning curves, and accessibility barriers for newcomers and academic researchers. Therefore, there remains a pressing need for an RLHF and RLVR framework that balances high performance, scalability, and ease of use—one that is straightforward enough for researchers new to the field, yet adaptable to diverse and evolving workloads.
In this paper, we present OpenRLHF, a simple, high-performance, fully open-source framework supporting both RLHF and RLVR, built upon Ray Liang et al. (2018), vLLM, DeepSpeed Yao et al. (2023), and HuggingFace Transformers Wolf et al. (2020). OpenRLHF offers three key contributions:
@@
Together, these innovations position OpenRLHF as an accessible yet powerful framework suitable for both efficient experimentation in academic environments and practical deployment scenarios. Already adopted by leading institutions and companies—including CMU, MIT, Microsoft, and HKUST—OpenRLHF demonstrates broad applicability and impact across the research community (see Appendix A for details on the framework’s broad impact and adoption). We publicly release OpenRLHF to foster openness, accelerate research, and promote innovation within the RLHF and RLVR ecosystem.
@@ 2 Related Work
Reinforcement Learning from Human Feedback and Reinforcement Learning with Verifiable Rewards. Reinforcement Learning from Human Feedback (RLHF) has emerged as a pivotal paradigm for aligning large language models with human preferences Christiano et al. (2017); Stiennon et al. (2020). The foundational RLHF framework trains a reward model from human preference data and optimizes the language model using reinforcement learning algorithms such as PPO Schulman et al. (2017). This approach has been successfully deployed in prominent models including InstructGPT Ouyang et al. (2022), ChatGPT, and GPT-4 Achiam et al. (2023). Building upon RLHF, Reinforcement Learning with Verifiable Rewards (RLVR) leverages automatically verifiable signals from mathematical verification, code execution, or other automated evaluation mechanisms Guo et al. (2025); Shen et al. (2025); Cobbe et al. (2021). While PPO remains dominant for its stability Schulman et al. (2017); Ziegler et al. (2019), alternative approaches such as Direct Preference Optimization (DPO) Rafailov et al. (2023) have gained attention for computational efficiency. However, both RLHF and RLVR methods require substantial computational resources and sophisticated distributed training strategies. The computational challenges have motivated various optimization techniques, including efficient inference engines Kwon et al. (2023), memory-efficient training strategies Rajbhandari et al. (2020), and distributed orchestration frameworks Yao et al. (2023). Nevertheless, existing solutions often sacrifice either performance for simplicity or accessibility for optimization, creating a gap that OpenRLHF aims to address.
RLHF (or RLVR) Frameworks. The computational complexity of RLHF and RLVR training has driven specialized framework development. General-purpose RL frameworks Liang et al. (2018); Dhariwal et al. (2017); Shen et al. (2020) designed for small-scale networks fail to address LLM-specific challenges and typically employ multi-controller architectures with complex inter-process communication, creating steep learning curves. RLHF-specific frameworks face significant performance-accessibility trade-offs. Open-source solutions like TRL von Werra et al. (2020), DeepSpeed-Chat Yao et al. (2023), and ColossalChat Li et al. (2023) provide accessible implementations but often lack sophisticated orchestration capabilities and struggle with inference optimization. Industrial frameworks like Nemo-aligner Shen et al. (2024a), ChatLearn alibaba (2017), and Verl Sheng et al. (2024) offer advanced optimizations including 3D parallelism and memory management. However, they feature tightly coupled architectures requiring substantial engineering expertise and extensive infrastructure setup, creating accessibility barriers for academic researchers. These systems adopt static resource allocation paradigms, leading to suboptimal utilization and limited adaptability. Most frameworks inadequately address both inference bottlenecks and usability challenges simultaneously, forcing users to choose between high performance and ease of use.
@@ 3.1 Overview: Ray-based RLHF architecture
Figure 1: Overall architecture of OpenRLHF. The system assigns GPUs to two primary roles: rollout engines dedicated to response generation and actor engines responsible for computing log-probabilities and model training. OpenRLHF leverages Ray for distributed scheduling, integrates VLLM to achieve efficient response rollout with low GPU memory usage, and employs DeepSpeed-Zero for 3D parallelism (including tensor, data, and sequence parallelism) to enable efficient model training. The underlying models are instantiated with flexible Transformer architectures, making the system easy to extend and adapt for diverse research scenarios.
OpenRLHF is the first open-source, Ray-based RLHF architecture that assigns a batch of GPUs to distinct roles and manages both data flow and workflow among these roles using Ray’s scheduling capabilities. As illustrated in Figure 3, it defines two primary roles: the rollout engine, responsible for response generation to given prompts, and the ZeRO engine, which computes logprobs, reference policy logprobs, and handles model training (For a detailed design of the PPO workflow, refer to Appendix B, and for in-depth implementation tips, consult Blog Shen et al. (2024b).)
Additionally, we leverage VLLM to implement the rollout engine, enabling efficient response generation with minimal GPU memory usage. For model training, we adopt DeepSpeed, implementing 3D parallelism — involving automatic tensor parallelism, ZeRO/data parallelism, and sequence parallelism — to efficiently train both the actor and value models.
Core reason for ease of use: simplified model slicing, seamless integration, and flexible scheduling. The exchange of model weights between the rollout engine and the training engine is enabled by a flexible slicing and partitioning pipeline. Hugging Face Transformer Jain (2022) models are instantiated and trained using DeepSpeed ZeRO, AutoTP, and Ring-Attention Liu et al. (2023) model parallelism. These model slices are then efficiently transferred to VLLM through AutoTP and AutoPP, which dynamically partition the models into sub-modules. The Ray-based scheduling mechanism enables seamless switching between different model parallelism modes, such as hybrid engine and asynchronous training. This streamlined workflow significantly reduces complexity, making the system highly user-friendly and easy to extend. Compared with architectures such as DeepSpeed-Chat, Transformer Reinforcement Learning (TRL), or other mainstream frameworks, OpenRLHF supports asynchronous dataflow and remote engine interactions, significantly improving the overall efficiency of the training process and the agent workflow.
@@ 3.2 Distributed and Efficient System Design
3D Parallelism with DeepSpeed-ZeRO and Ring Attention. To enable seamless and efficient scalability for large models, OpenRLHF integrates the latest automatic tensor parallelism (AutoTP) feature from DeepSpeed-ZeRO. In many industrial-grade RLHF architectures, users previously needed to manually specify an injection policy for each transformer model, identifying the linear layers and attention outputs that required communication between data-parallel ranks. In contrast, OpenRLHF leverages DeepSpeed-ZeRO’s new capability to support automatic tensor parallelism for HuggingFace models by default: when kernel injection is not enabled and an injection policy is not provided, DeepSpeed automatically determines and applies the necessary policy at runtime. This greatly simplifies the user experience and extends robust tensor parallelism support to a wider range of models, removing the need for complex engineering or manual configuration.
@@
In addition, OpenRLHF implements sequence parallelism through ring attention. Ring attention employs a ring-based communication topology, efficiently distributing attention computation for long sequences across multiple GPUs while minimizing both memory usage and communication overhead. This is especially critical for modern RLHF and RLVR workloads involving long chain-of-thought (CoT) reasoning, where conventional attention computation can become a major scalability bottleneck. By combining AutoTP, data parallelism, and ring attention-based sequence parallelism, OpenRLHF empowers large-scale, efficient, and highly usable RLHF model training on flexible GPU clusters.
@@
Accelerated CoT Inference with vLLM. As LLMs advance in reasoning, RLHF and RLVR pipelines increasingly face inference bottlenecks, especially with long chain-of-thought (CoT) outputs. For models like OpenAI-o1 and DeepSeek-R1, CoT generation can dominate training time, making efficient long-form inference crucial for scalability. To address this challenge, OpenRLHF integrates the vLLM inference framework, which is specifically designed for high-throughput and memory-efficient LLM serving. vLLM provides a streamlined interface for generating RLHF samples and supporting frequent model weight updates.
@@
The core innovation in vLLM is efficient management of attention key and value memory with PagedAttention Kwon et al. (2023). This technique significantly reduces memory waste to less than 4%, enabling the batching of more sequences and enhancing GPU utilization and throughput. PagedAttention also supports efficient memory sharing for advanced sampling methods, such as parallel sampling and beam search, reducing memory usage by up to 55% and further boosting inference efficiency. Besides PagedAttention, vLLM has several other advantages, including continuous batching of incoming requests, fast model execution with CUDA Graph, and CUDA kernels optimized with FlashAttention and FlashInference, which enable rapid and memory-efficient attention computations. It also features speculative decoding for faster inference and chunked prefill to reduce latency on long sequences. Collectively, these enhancements make vLLM highly effective for large-scale, long-sequence inference, especially suitable for RLHF and RLVR pipelines, where efficiency and scalability are crucial.
@@
Asynchronous Dataflow and Remote Engine Interactions. OpenRLHF supports asynchronous dataflow and remote engine communication to maximize system throughput and resource utilization during distributed RLHF training. In this architecture, rollout engines, actor engines, and remote engines operate independently and communicate via message passing, enabling immediate processing as soon as data becomes available. Fully asynchronous execution is especially important in the chain-of-thought (CoT) era, where inference involves generating multi-step reasoning that can vary greatly in length and computational cost. In synchronous frameworks, the slowest CoT generation can bottleneck the whole pipeline and waste resources. In contrast, OpenRLHF’s asynchronous design allows each engine to operate at its own pace, ensuring hardware is utilized even for long or variable CoT tasks. This not only accelerates training and evaluation, but also enables efficient scaling and supports dynamic agent RL workflows. Leveraging asynchronous remote engine interactions, OpenRLHF is readily extensible for scalable agent RL training in modern, CoT-centric environments.
@@
OpenRLHF Verl OpenRLHF Verl OpenRLHF Verl OpenRLHF Verl

### v6

@@ Abstract
Large Language Models (LLMs) fine-tuned via Reinforcement Learning from Human Feedback (RLHF) and Reinforcement Learning with Verifiable Rewards (RLVR) significantly improve the alignment of human-AI values, further raising the upper bound of AI capabilities, particularly in reasoning-intensive, long-context Chain-of-Thought (CoT) tasks. However, existing frameworks commonly face challenges such as inference bottlenecks and complexity barriers, which restrict their accessibility to newcomers. To bridge this gap, we introduce OpenRLHF, a user-friendly, scalable, and easy-to-learn open-source RLHF framework built upon Ray, vLLM, DeepSpeed, and HuggingFace Transformers, featuring a simplified design, clear code structure, and comprehensive documentation to facilitate entry for researchers and practitioners. Experimental results show that OpenRLHF achieves superior training efficiency, with speedups ranging from 1.22× to 1.68× across different model sizes, compared to state-of-the-art frameworks. Additionally, it requires significantly fewer lines of code for implementation. OpenRLHF is publicly available at https://github.com/OpenRLHF/OpenRLHF, and has already been adopted by leading institutions to accelerate RLHF research and learning.
@@ 1 Introduction
Large Language Models (LLMs) fine-tuned via Reinforcement Learning from Human Feedback (RLHF) and Reinforcement Learning with Verifiable Rewards (RLVR) have markedly advanced human-AI alignment and elevated the upper bound of AI capabilities [1, 2, 3, 4]. These approaches enable models to better conform to human intentions and values while achieving superior reasoning performance. Notably, models such as GPT-4 [5], DeepSeek-R1 [3], and Claude [6] excel at complex reasoning tasks by generating detailed step-by-step rationales, commonly referred to as Chain-of-Thought (CoT) outputs.
However, RLHF and RLVR training methodologies—especially those employing Proximal Policy Optimization (PPO)—face significant computational challenges. In particular, the inference phase often accounts for over 90% of the total RLHF (or RLVR) runtime, as models need to generate thousands of tokens during each inference step. Consequently, there is an increasing demand for efficient and scalable frameworks that reduce inference overhead and simplify the training workflows for distributed RLHF and RLVR.
Existing RLHF and RLVR systems, such as DeepSpeed-Chat [7], TRL [8], and ColossalChat [9], have made notable progress in distributed computation and memory efficiency. Conversely, industrial-grade solutions like Nemo-aligner [10] and ChatLearn [11] offer advanced optimizations at the modeling and framework levels. While verl [12], a framework proposed after our initial development, also provides sophisticated optimizations (e.g., its 3D-Hybrid engine), these industrial solutions generally feature tightly coupled and specialized designs that introduce considerable complexity, steep learning curves, and accessibility barriers for newcomers and academic researchers. Yet, their tightly coupled and specialized designs introduce significant complexity, steep learning curves, and accessibility barriers for newcomers and academic researchers. Therefore, there remains a pressing need for an RLHF and RLVR framework that balances high performance, scalability, and ease of use—one that is straightforward enough for researchers new to the field, yet adaptable to diverse and evolving workloads.
In this paper, we present OpenRLHF, a simple, high-performance, fully open-source framework supporting both RLHF and RLVR, built upon Ray [13], vLLM, DeepSpeed [7], and HuggingFace Transformers [14]. OpenRLHF offers these key contributions:
@@
Together, these innovations position OpenRLHF as an accessible yet robust framework suitable for both efficient experimentation in academic environments and practical deployment scenarios. Already adopted by leading institutions and companies—including CMU, MIT, Microsoft, and HKUST—OpenRLHF demonstrates broad applicability and impact across the research community (see Appendix B for details on the framework’s broad impact and adoption). We publicly release OpenRLHF to foster openness, accelerate research, and promote innovation within the RLHF and RLVR ecosystem.
@@ 2 Related Work
RLHF and RLVR
Reinforcement Learning from Human Feedback (RLHF) has emerged as a pivotal paradigm for aligning large language models with human preferences [1, 2]. The foundational RLHF framework trains a reward model using human preference data and optimizes the language model with reinforcement learning algorithms, such as PPO [15]. This approach has been successfully deployed in prominent models, including InstructGPT [16], ChatGPT, and GPT-4 [5]. Building upon RLHF, Reinforcement Learning with Verifiable Rewards (RLVR) leverages automatically verifiable signals from mathematical verification, code execution, or other automated evaluation mechanisms [3, 4, 17]. While PPO remains dominant due to its stability [15, 18], alternative approaches, such as Direct Preference Optimization (DPO) [19], have gained attention for their computational efficiency. However, both RLHF and RLVR methods require substantial computational resources and sophisticated distributed training strategies. The computational challenges have motivated various optimization techniques, including efficient inference engines [20], memory-efficient training strategies [21], and distributed orchestration frameworks [7]. Nevertheless, existing solutions often sacrifice either performance for simplicity or accessibility for optimization, creating a gap that OpenRLHF aims to address.
RLHF (RLVR) Frameworks
The computational complexity of RLHF and RLVR training has driven the development of specialized frameworks for these tasks. General-purpose RL frameworks [13, 22] designed for small-scale networks often fail to address LLM-specific challenges and typically employ multi-controller architectures with complex inter-process communication, resulting in steep learning curves. RLHF-specific frameworks face significant performance-accessibility trade-offs. Open-source solutions like TRL [8], DeepSpeed-Chat [7], and ColossalChat [9] provide accessible implementations but often lack sophisticated orchestration capabilities and struggle with inference optimization. Industrial frameworks like Nemo-aligner [10], ChatLearn [11], and verl [12] offer advanced optimizations including 3D parallelism and memory management. However, they feature tightly coupled architectures requiring substantial engineering expertise and extensive infrastructure setup, creating accessibility barriers for academic researchers. These systems employ static resource allocation paradigms, resulting in suboptimal utilization and limited adaptability. Most frameworks inadequately address both inference bottlenecks and usability challenges simultaneously, forcing users to choose between high performance and ease of use.
@@ 3.1 Overview: Ray-based RLHF architecture
Figure 1: Overall architecture of OpenRLHF. The system assigns GPUs to two primary roles: rollout engines dedicated to response generation and actor engines responsible for computing log-probabilities and model training. OpenRLHF leverages Ray for distributed scheduling, integrates vLLM to achieve efficient response rollout with low GPU memory usage, and employs DeepSpeed ZeRO for 3D parallelism (including tensor, data, and sequence parallelism) to enable efficient training. The underlying models are instantiated with flexible Transformer architectures, making the system easy to extend and adapt for diverse scenarios.
OpenRLHF is the first open-source, Ray-based RLHF architecture that assigns a batch of GPUs to distinct roles and manages both data flow and workflow among these roles using Ray’s scheduling capabilities. As illustrated in Figure 1, it defines two primary roles: the rollout engine, responsible for response generation to given prompts, and the ZeRO engine, which computes logprobs, reference policy logprobs, and handles model training (For a detailed design of the PPO workflow, refer to Appendix C, and for in-depth implementation tips, consult Blog [23].)
Additionally, we leverage vLLM as the rollout engine, enabling efficient response generation with minimal GPU memory usage. For model training, we adopt DeepSpeed, which implements 3D parallelism, including automatic tensor parallelism, ZeRO/data parallelism, and sequence parallelism, to train both the actor and value models efficiently.
Core reason for ease of use
The remarkable usability of our framework stems from three core design principles: simplified model slicing, seamless integration, and flexible scheduling, which collectively streamline the workflow and significantly reduce the implementation burden during training and deployment. The exchange of model weights between the rollout engine and the training engine is enabled by a flexible slicing and partitioning pipeline. Hugging Face Transformer [24] models are instantiated and trained using DeepSpeed ZeRO, AutoTP, and Ring-Attention [25] model parallelism. These model slices are then efficiently transferred to vLLM through AutoTP and AutoPP, which dynamically partition the models into sub-modules. The Ray-based scheduling mechanism enables seamless switching between different model parallelism modes, such as hybrid engine and asynchronous training. This streamlined workflow significantly reduces complexity, making the system highly user-friendly and easy to extend. Compared with architectures such as DeepSpeed-Chat, Transformer Reinforcement Learning (TRL), or other mainstream frameworks, OpenRLHF supports asynchronous dataflow and remote engine interactions, significantly improving the overall efficiency of the training process and the agent workflow.
@@ 3.2 Distributed and Efficient System Design
3D Parallelism with DeepSpeed ZeRO and Ring Attention
@@
To enable seamless and efficient scalability for large models, OpenRLHF integrates the latest automatic tensor parallelism (AutoTP) feature from DeepSpeed ZeRO. In many industrial-grade RLHF architectures, users previously needed to manually specify an injection policy for each transformer model, identifying the linear layers and attention outputs that required communication between data-parallel ranks. In contrast, OpenRLHF leverages DeepSpeed ZeRO’s new capability to support automatic tensor parallelism for HuggingFace models by default. When kernel injection is not enabled and an injection policy is not provided, DeepSpeed automatically determines and applies the necessary policy at runtime. It can dramatically simplify the user experience and extend robust tensor parallelism support to a broader range of models, removing the need for complex engineering or manual configuration.
@@
In addition, OpenRLHF implements sequence parallelism through ring attention. Ring attention employs a ring-based communication topology, efficiently distributing attention computation for long sequences across multiple GPUs while minimizing both memory usage and communication overhead. It is especially critical for modern RLHF and RLVR workloads involving long CoT reasoning, where conventional attention computation can become a major scalability bottleneck. By combining AutoTP, data parallelism, and ring attention-based sequence parallelism, OpenRLHF empowers large-scale, efficient, and highly usable RLHF model training on flexible GPU clusters.
Accelerated CoT Inference with vLLM
@@
As LLMs advance in reasoning, RLHF and RLVR pipelines increasingly face inference bottlenecks, particularly with long CoT outputs. For models like OpenAI-o1 and DeepSeek-R1, CoT generation can dominate training time, making efficient long-form inference crucial for scalability. To address this challenge, OpenRLHF integrates the vLLM inference framework, which is specifically designed for high-throughput and memory-efficient LLM serving. vLLM provides a streamlined interface for generating RLHF samples and supporting frequent model weight updates.
@@
The core innovation in vLLM is efficient management of attention key and value memory with PagedAttention [20]. The technique significantly reduces memory waste to less than 4%, enabling the batching of more sequences and enhancing GPU utilization and throughput. PagedAttention also supports efficient memory sharing for advanced sampling methods, such as parallel sampling and beam search, reducing memory usage by up to 55% and further boosting inference efficiency. Besides PagedAttention, vLLM has several other advantages, including continuous batching of incoming requests, fast model execution with CUDA Graph, and CUDA kernels optimized with FlashAttention and FlashInference, which enable rapid and memory-efficient attention computations. It also features speculative decoding for faster inference and chunked prefill to reduce latency on long sequences. Collectively, these enhancements make vLLM highly effective for large-scale, long-sequence inference, especially suitable for RLHF and RLVR pipelines, where efficiency and scalability are crucial.
Asynchronous Dataflow and Remote Engine Interactions
@@
OpenRLHF supports asynchronous dataflow and remote engine communication to maximize system throughput and resource utilization during distributed RLHF training. In this architecture, rollout engines, actor engines, and remote engines operate independently and communicate via message passing, enabling immediate processing as soon as data becomes available. Fully asynchronous execution is critical in the CoT era, where inference involves generating multi-step reasoning that can vary significantly in length and computational cost. In synchronous frameworks, the slowest CoT generation can bottleneck the whole pipeline and waste resources. In contrast, OpenRLHF’s asynchronous design allows each engine to operate at its own pace, ensuring hardware is utilized even for long or variable CoT tasks. It not only accelerates training and evaluation but also enables efficient scaling and supports dynamic agent RL workflows. Leveraging asynchronous remote engine interactions, OpenRLHF is readily extensible for scalable agent RL training in modern, CoT-centric environments.
@@
Ours verl Ours verl Ours verl Ours verl